OpenStudy (anonymous):

what are the coordinates of the y-intercept of the line -x-7y=21?

OpenStudy (anonymous):

For all y-intercepts, x = 0. All we have to do is plug 0 in the equation for x and solve for y: 0 - 7y = 21 -7y = 21 y = -3 So your y-intercept is (0, -3).